# Correlations Are Ruining Your Gradient Descent

### v2

@@ 2.1 Gradient Descent
where, ${\delta\bm{\theta}}^{\top}\delta\bm{\theta}=||\delta\bm{\theta}||^{2}_{2}=\sum_{i}\delta\bm{\theta}_{i}^{2}$. Note, one can derive gradient descent by simply finding the minimum of this optimisation function (where the derivative with respect to $\delta\bm{\theta}$ is zero). Thus, gradient descent assumes that it is sensible to measure and limit the distance of our parameter update in terms of the squared (Euclidean) norm of the parameter change. Natural gradients supposes that this is not the best choice.
@@ 2.3 Natural Gradient Descent for a DNN
$\bm{G}_{\bm{W}_{i}}=\langle\nabla_{\theta_{\bm{W}_{i}}}\ell^{\top}\,\nabla_{\bm{\theta}_{\bm{W}_{i}}}\ell\rangle_{\bm{x}_{0}}=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)^{\top}\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)\right\rangle_{\bm{x}_{0}}\$
@@
${\delta\bm{\theta}}_{\bm{W}_{i}}=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}^{\top}\frac{\partial\ell}{\partial\bm{h}_{i}}\right\rangle^{-1}_{\bm{y}}\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right\rangle_{\bm{x}_{i-1}}\left\langle\bm{x}_{i-1}^{\top}\bm{x}_{i-1}\right\rangle^{-1}_{\bm{x}_{i-1}}\right).\$
@@ 3 Simple decorrelation mechanisms can rid networks of data correlations
where $\eta_{M}$ is a learning rate for this update, and $g_{i}$ is a scalar gain which ensures that each of the layer’s decorrelated outputs has the same norm as it had prior to decorrelation, $g_{i}=||\bm{x}_{i}^{2}||_{F}/||\bar{\bm{x}}_{i}^{2}||_{F}$. This update minimises a loss capturing the correlations in $\bar{\bm{x}}$ in the style of Ahmad et al. [2022]. Updates can be carried out in a stochastic fashion by simply measuring the correlations in $\bar{x}$ within a minibatch for iterative updating, see Appendix D for the full pseudocode for such updating. Note that one can further improve upon the level of decorrelation by de-meaning the input data prior to this decorrelation process, such that $\hat{\bm{x}_{i}}=\bm{M}_{i}(\bm{x}_{i}-\bm{\mu}_{i})$ where $\bm{\mu}_{i}$ is a unit-wise learned mean or batch-wise computed mean. We find this to further improve performance in practice.
@@
Finally, aside from each of these benefits, this system can be easily converted to a system of recurrent dynamics with local-information, see Appendix C. This is particularly a benefit if one wished to implement this method in a neuromorphic device or to link it to biological nervous systems.
@@ 4 Approximate methods for gradient descent work significantly better when coupled with decorrelation
Figure 4: Adding a decorrelation mechanism at every layer of a neural network can massively speed up training convergence speed. Results are shown for backpropagation, feedback alignment, and node perturbation train upon the CIFAR10 in four hidden-layer dense networks, and for backpropagation and feedback alignment upon the CIFAR100 and TinyImageNet classification tasks in five hidden-layer convolutional neural networks. All learning methods were combined with the Adam optimizer [Kingma and Ba, 2014] and the categorical cross entropy loss. Network architectures and training hyperparameters are described in detail in Appendix E. Envelopes show the max and min accuracy levels across five randomly seeded networks. Parameters were selected based upon an independent grid search over decorrelation learning rates and forward learning rates for maximum validation set performance.
@@
Here we show that not only does decorrelation improve training of multi-layer neural networks with the node-perturbation algorithm (as has already been shown by Dalm et al. [2023]) but also significantly improves multi-layer network training when combined with implementation of the feedback alignment algorithm. Networks parameters in pseudocode, as well as the hyperparameters which were used for training are presented in Appendix E and an example of the computational pseudocode provided in Appendix D. Code to reproduce these simulations is available at https://github.com/nasiryahm/CorrelationsRuinGD. Note that all simulations shown below are presented based upon a parameter grid search which was carried out individually for each credit assignment method and best parameters selected for each curve based upon a validation set.
@@ Appendix A Loss distance the approriate Riemann metric for natural gradients
$(\ell(f(\theta+\delta\theta,x_{0}),y)-\ell(f(\theta,x_{0}),y))^{2}$ $=(\ell+\nabla_{\theta}\ell{\delta\theta}+{\delta\theta}^{\top}\nabla^{2}_{\theta}\ell{\delta\theta}+...-\;\ell)^{2}$ (1)
$=(\nabla_{\theta}\ell{\delta\theta}+{\delta\theta}\nabla^{2}_{\theta}\ell{\delta\theta}+...)^{2}$ (2)
$={\delta\theta}^{\top}\nabla_{\theta}\ell^{\top}\nabla_{\theta}\ell{\delta\theta}+2(\nabla_{\theta}\ell{\delta\theta}\times{\delta\theta}^{\top}\nabla^{2}_{\theta}\ell{\delta\theta})+...$ (3)
@@
${\delta\theta}^{\top}G\delta\theta={\delta\theta}^{\top}\left\langle\nabla_{\theta}\ell^{\top}\nabla_{\theta}\ell\right\rangle_{x_{0}\in\mathcal{D}}{\delta\theta}.$
@@ Appendix B Derivation for the Riemann metric for a deep neural network
$G(\theta)$ $=\left\langle\nabla_{\theta}\ell^{\top}\nabla_{\theta}\ell\right\rangle_{x_{0}}$
$=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial h_{1}}x_{0}^{\top},\frac{\partial\ell}{\partial h_{2}}x_{1}^{\top},...\frac{\partial\ell}{\partial h_{L}}x_{L-1}^{\top}\right)^{\top}\text{Vec}\left(\frac{\partial\ell}{\partial h_{1}}x_{0}^{\top},\frac{\partial\ell}{\partial h_{2}}x_{1}^{\top},...\frac{\partial\ell}{\partial h_{L}}x_{L-1}^{\top}\right)\right\rangle_{x_{0}}$
@@
$G_{ij}=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial h_{j}}x_{j-1}^{\top}\right)^{\top}\text{Vec}\left(\frac{\partial\ell}{\partial h_{i}}x_{i-1}^{\top}\right)\right\rangle_{x_{0}}\\ =\left\langle\left(x_{j-1}\otimes\frac{\partial\ell}{\partial h_{j}}\right)^{\top}\left(x_{i-1}\otimes\frac{\partial\ell}{\partial h_{i}}\right)\right\rangle_{x_{0}}\$
@@
$G_{ij}=\left\langle\left(x_{j-1}^{\top}x_{i-1}\right)\otimes\left(\frac{\partial\ell}{\partial h_{j}}^{\top}\frac{\partial\ell}{\partial h_{i}}\right)\right\rangle_{x_{0}}=\left\langle x_{j-1}^{\top}x_{i-1}\right\rangle_{x_{0}}\otimes\left\langle\frac{\partial\ell}{\partial h_{j}}^{\top}\frac{\partial\ell}{\partial h_{i}}\right\rangle_{y}.$
@@
${\delta\theta}_{W_{i}}$ $=-{\eta}\langle\nabla_{\theta_{W_{i}}}\ell^{\top}\,\nabla_{\theta_{W_{i}}}\ell\rangle_{x_{0}}^{-1}\nabla_{\theta_{W_{i}}}\mathcal{L}$ (5)
$=-{\eta}\left(\left\langle x_{i-1}^{\top}x_{i-1}\right\rangle_{x_{0}}\otimes\left\langle\frac{\partial\ell}{\partial h_{i}}^{\top}\frac{\partial\ell}{\partial h_{i}}\right\rangle_{y}\right)^{-1}\nabla_{\theta_{W_{i}}}\mathcal{L}$ (6)
$=-{\eta}\left(\left\langle x_{i-1}^{\top}x_{i-1}\right\rangle^{-1}_{x_{0}}\otimes\left\langle\frac{\partial\ell}{\partial h_{i}}^{\top}\frac{\partial\ell}{\partial h_{i}}\right\rangle^{-1}_{y}\right)\text{Vec}\left\langle\frac{\partial\ell}{\partial h_{i}}x_{i-1}^{\top}\right\rangle_{x_{0}}$ (7)
$=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial h_{i}}^{\top}\frac{\partial\ell}{\partial h_{i}}\right\rangle^{-1}_{y}\left\langle\frac{\partial\ell}{\partial h_{i}}x_{i-1}^{\top}\right\rangle_{x_{0}}\left\langle x_{i-1}^{\top}x_{i-1}\right\rangle^{-1}_{x_{0}}\right)$ (8)
@@
Appendix C Our decorrelation mechanism as a recurrent system of dynamics
One may describe our proposed decorrelation mechanism as a system of lateral dynamics where $\frac{d\bar{\bm{x}}}{dt}=\bar{\bm{x}}-{\bm{M}}^{-1}\bm{x}$. This system can be shown to precisely arrive at the exact solution as outlined in the above text, $\bar{\bm{x}}=\bm{M}\bm{x}$. Furthermore, if one wished to update the parameters of this dynamic decorrelation setup in stochastic manner (single sample mini-batches), then one may apply the Sherman-Morrison formula to show that $\bm{M}^{-1}$ should be updated with
$\bm{M}^{-1}\leftarrow(g(\bm{I}-\eta_{M}\bar{\bm{x}}\bar{\bm{x}}^{\top})\bm{M})^{-1}$ $=g^{-1}\bm{M}^{-1}(I+\frac{\eta_{M}\bar{\bm{x}}\bar{\bm{x}}^{\top}}{1-\eta_{M}\bar{\bm{x}}^{\top}\bar{\bm{x}}})$
$\approx{g}^{-1}\bm{M}^{-1}(I+\eta_{M}\bar{\bm{x}}\bar{\bm{x}}^{\top})$
$\approx{g}^{-1}(\bm{M}^{-1}+\eta_{M}(\bm{M}^{-1}\bar{\bm{x}})\bar{\bm{x}}^{\top}).$
Assuming that the total decorrelating signal to each individual unit can be locally measured ($\bm{M}^{-1}\bar{\bm{x}}$) then this system of dynamics can also be updated in a local fashion by each node.
Appendix D Training pseudocode
@@
     $\hat{\bm{x}}_{i}=\bm{M}_{i}(\bm{x}_{i}-\bm{\mu}_{i})$triangleright Demean and decorrelate state at layer $i$
     $\bm{x}_{i+1}=\phi(\bm{W}_{i+1}\hat{\bm{x}}_{i})$triangleright Pass forward to the next layer
@@
         $\bm{\delta}_{L}=\phi^{\prime}(\bm{x}_{L})(\bm{x}_{L}-\bm{y})$triangleright Compute gradient at output layer (regression or CCE)
@@
         $\bm{\delta}_{i}=\phi^{\prime}(\bm{x}_{i})\bm{M}_{i}^{\top}\bm{W}_{i+1}^{\top}\bm{\delta}_{i+1}$triangleright Backpropagate (or replace $\bm{W}_{i+1}$ for FA)
@@
     $\bm{C}=\hat{\bm{x}}_{i}\hat{\bm{x}}_{i}^{\top}$triangleright Compute remaining correlation
     $\bm{g}=\sqrt{\bm{x}_{i}^{2}/\hat{\bm{x}}_{i}^{2}}$triangleright Compute rescaling
     $\bm{M}_{i}\leftarrow\bm{g}\circ(\bm{M}_{i}-\eta_{M}\bm{C}\bm{M}_{i})$triangleright Update decorrelation matrix
     $\bm{\mu}_{i}\leftarrow\bm{\mu}_{i}+0.1(\bm{x}_{i}-\bm{\mu}_{i})$triangleright Update mean estimate (fixed learning rate)
@@
     $\bm{W}_{i+1}\leftarrow\bm{W}_{i+1}-\eta_{W}\bm{\delta}_{i+1}\hat{\bm{x}}_{i}^{\top}$triangleright Update by SGD (alternatively by other optimizer)
@@
Appendix E Architectures and Hyperparameters

### v3

@@ 2.1 Gradient Descent
[…]
@@ 2.3 Natural Gradient Descent for a DNN
$\bm{G}_{\bm{W}_{i}}=\langle\nabla_{\theta_{\bm{W}_{i}}}\ell^{\top}\,\nabla_{\bm{\theta}_{\bm{W}_{i}}}\ell\rangle_{\bm{x}_{0}}=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)^{\top}\text{Vec}\left(\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right)\right\rangle_{\bm{x}_{0}}\\$
@@
${\delta\bm{\theta}}_{\bm{W}_{i}}=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}^{\top}\frac{\partial\ell}{\partial\bm{h}_{i}}\right\rangle^{-1}_{\bm{y}}\left\langle\frac{\partial\ell}{\partial\bm{h}_{i}}\bm{x}_{i-1}^{\top}\right\rangle_{\bm{x}_{i-1}}\left\langle\bm{x}_{i-1}^{\top}\bm{x}_{i-1}\right\rangle^{-1}_{\bm{x}_{i-1}}\right).\\$
@@ 3 Simple decorrelation mechanisms can rid networks of data correlations
[…]
@@
Finally, aside from each of these benefits, this system can be easily converted to a system of recurrent dynamics with local-information, see Appendix D. This is particularly a benefit if one wished to implement this method in a neuromorphic device or to link it to biological nervous systems.
@@ 4 Approximate methods for gradient descent work significantly better when coupled with decorrelation
Figure 4: Adding a decorrelation mechanism at every layer of a neural network can massively speed up training convergence speed. Results are shown for backpropagation, feedback alignment, and node perturbation train upon the CIFAR10 in four hidden-layer dense networks, and for backpropagation and feedback alignment upon the CIFAR100 and TinyImageNet classification tasks in five hidden-layer convolutional neural networks. All learning methods were combined with the Adam optimizer [Kingma and Ba, 2014] and the categorical cross entropy loss. Network architectures and training hyperparameters are described in detail in Appendix F. Envelopes show the max and min accuracy levels across five randomly seeded networks. Parameters were selected based upon an independent grid search over decorrelation learning rates and forward learning rates for maximum validation set performance.
@@
Here we show that not only does decorrelation improve training of multi-layer neural networks with the node-perturbation algorithm (as has already been shown by Dalm et al. [2023]) but also significantly improves multi-layer network training when combined with implementation of the feedback alignment algorithm. Networks parameters in pseudocode, as well as the hyperparameters which were used for training are presented in Appendix F and an example of the computational pseudocode provided in Appendix E. Code to reproduce these simulations is available at https://github.com/nasiryahm/CorrelationsRuinGD. Note that all simulations shown below are presented based upon a parameter grid search which was carried out individually for each credit assignment method and best parameters selected for each curve based upon a validation set.
@@ Appendix A Loss distance the approriate Riemann metric for natural gradients
$(\ell(f(\theta+\delta\theta,x_{0}),y)-\ell(f(\theta,x_{0}),y))^{2}$ $=(\ell+\nabla_{\theta}\ell\cdot{\delta\theta}+{\delta\theta}^{\top}\nabla^{2}_{\theta}\ell{\delta\theta}+...-\;\ell)^{2}$ (1)
$=({\delta\theta}^{\top}\nabla_{\theta}\ell+{\delta\theta}\nabla^{2}_{\theta}\ell{\delta\theta}+...)^{2}$ (2)
$={\delta\theta}^{\top}\nabla_{\theta}\ell\nabla_{\theta}\ell^{\top}{\delta\theta}+2(\nabla_{\theta}\ell{\delta\theta}\times{\delta\theta}^{\top}\nabla^{2}_{\theta}\ell{\delta\theta})+$... (3)
@@
${\delta\theta}^{\top}G\delta\theta={\delta\theta}^{\top}\left\langle\nabla_{\theta}\ell\nabla_{\theta}\ell^{\top}\right\rangle_{x_{0}\in\mathcal{D}}{\delta\theta}.$
@@ Appendix B Derivation for the Riemann metric for a deep neural network
$G(\theta)$ $=\left\langle\nabla_{\theta}\ell\nabla_{\theta}\ell^{\top}\right\rangle_{x_{0}}$
$=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial h_{1}}x_{0}^{\top},\frac{\partial\ell}{\partial h_{2}}x_{1}^{\top},...\frac{\partial\ell}{\partial h_{L}}x_{L-1}^{\top}\right)\text{Vec}\left(\frac{\partial\ell}{\partial h_{1}}x_{0}^{\top},\frac{\partial\ell}{\partial h_{2}}x_{1}^{\top},...\frac{\partial\ell}{\partial h_{L}}x_{L-1}^{\top}\right)^{\top}\right\rangle_{x_{0}}$
@@
$G_{ij}=\left\langle\text{Vec}\left(\frac{\partial\ell}{\partial h_{j}}x_{j-1}^{\top}\right)\text{Vec}\left(\frac{\partial\ell}{\partial h_{i}}x_{i-1}^{\top}\right)^{\top}\right\rangle_{x_{0}}\\ =\left\langle\left(x_{j-1}\otimes\frac{\partial\ell}{\partial h_{j}}\right)\left(x_{i-1}\otimes\frac{\partial\ell}{\partial h_{i}}\right)^{\top}\right\rangle_{x_{0}}\\$
@@
$G_{ij}=\left\langle\left(x_{j-1}x_{i-1}^{\top}\right)\otimes\left(\frac{\partial\ell}{\partial h_{j}}\frac{\partial\ell}{\partial h_{i}}^{\top}\right)\right\rangle_{x_{0}}=\left\langle x_{j-1}x_{i-1}^{\top}\right\rangle_{x_{0}}\otimes\left\langle\frac{\partial\ell}{\partial h_{j}}\frac{\partial\ell}{\partial h_{i}}^{\top}\right\rangle_{y}.$
@@
${\delta\theta}_{W_{i}}$ $=-{\eta}\langle\nabla_{\theta_{W_{i}}}\ell\,\nabla_{\theta_{W_{i}}}\ell^{\top}\rangle_{x_{0}}^{-1}\nabla_{\theta_{W_{i}}}\mathcal{L}$ (5)
$=-{\eta}\left(\left\langle x_{i-1}x_{i-1}^{\top}\right\rangle_{x_{0}}\otimes\left\langle\frac{\partial\ell}{\partial h_{i}}\frac{\partial\ell}{\partial h_{i}}^{\top}\right\rangle_{y}\right)^{-1}\nabla_{\theta_{W_{i}}}\mathcal{L}$ (6)
$=-{\eta}\left(\left\langle x_{i-1}x_{i-1}^{\top}\right\rangle^{-1}_{x_{0}}\otimes\left\langle\frac{\partial\ell}{\partial h_{i}}\frac{\partial\ell}{\partial h_{i}}^{\top}\right\rangle^{-1}_{y}\right)\text{Vec}\left\langle\frac{\partial\ell}{\partial h_{i}}x_{i-1}^{\top}\right\rangle_{x_{0}}$ (7)
$=-{\eta}\text{Vec}\left(\left\langle\frac{\partial\ell}{\partial h_{i}}\frac{\partial\ell}{\partial h_{i}}^{\top}\right\rangle^{-1}_{y}\left\langle\frac{\partial\ell}{\partial h_{i}}x_{i-1}^{\top}\right\rangle_{x_{0}}\left\langle x_{i-1}x_{i-1}^{\top}\right\rangle^{-1}_{x_{0}}\right)$ (8)
@@
Appendix C A note on decorrelation vs whitening
Wadia et al. [2021] pointed out that the restriction to whitened data can restrict the space of generalization unless properly regularized. This issue is one to keep in mind when developing methods which speed up training of models significantly and generalization performance should be a point of concern. However, as also described by Wadia et al. [2021], it may be that a regularized form of whitening/decorrelation would in fact be optimal for training for maximum generalization performance. Ultimately, we opt to avoid enforcing strict whitening and rely on decorrelation as an alternative and find it to be performative in practice.
Appendix D Our decorrelation mechanism as a recurrent system of dynamics
One may describe our proposed decorrelation mechanism as a system of lateral dynamics where $\frac{d\bar{\bm{x}}}{dt}=\bar{\bm{x}}-\bm{R}\bm{x}$, where $\bm{R}={\bm{M}}^{-1}$. This system can be shown to precisely arrive at the exact solution as outlined in the above text, $\bar{\bm{x}}=\bm{M}\bm{x}$. Furthermore, if one wished to update the parameters of this dynamic decorrelation setup in stochastic manner (single sample mini-batches), then one may apply the Sherman-Morrison formula to show that $\bm{M}^{-1}$ should be updated with
$\bm{M}^{-1}\leftarrow(\bm{g}\circ(\bm{I}-\eta_{M}\bar{\bm{x}}\bar{\bm{x}}^{\top})\bm{M})^{-1}$ $=\bm{M}^{-1}(I+\frac{\eta_{M}\bar{\bm{x}}\bar{\bm{x}}^{\top}}{1-\eta_{M}\bar{\bm{x}}^{\top}\bar{\bm{x}}})\circ\bm{g}^{-1}$
$\approx\bm{M}^{-1}(I+\eta_{M}\bar{\bm{x}}\bar{\bm{x}}^{\top})\circ\bm{g}^{-1}$
$\approx(\bm{M}^{-1}+\eta_{M}(\bm{M}^{-1}\bar{\bm{x}})\bar{\bm{x}}^{\top})\circ\bm{g}^{-1}.$
After conversion to the notation with matrix, $\bm{R}$,
$\bm{R}\leftarrow(\bm{R}+\eta_{M}(\bm{R}\bar{\bm{x}})\bar{\bm{x}}^{\top})\circ\bm{g}^{-1}.$
Assuming that the total decorrelating signal to each individual unit can be locally measured ($\bm{R}\bar{\bm{x}}$) then this system of dynamics can also be updated in a local fashion by each node.
Appendix E Training pseudocode
@@
  $\hat{\bm{x}}_{i}=\bm{M}_{i}(\bm{x}_{i}-\bm{\mu}_{i})$ $\triangleright$ Demean and decorrelate state at layer $i$
  $\bm{x}_{i+1}=\phi(\bm{W}_{i+1}\hat{\bm{x}}_{i})$ $\triangleright$ Pass forward to the next layer
@@
   $\bm{\delta}_{L}=\phi^{\prime}(\bm{x}_{L})(\bm{x}_{L}-\bm{y})$ $\triangleright$ Compute gradient at output layer (regression or CCE)
@@
   $\bm{\delta}_{i}=\phi^{\prime}(\bm{x}_{i})\bm{M}_{i}^{\top}\bm{W}_{i+1}^{\top}\bm{\delta}_{i+1}$ $\triangleright$ Backpropagate (or replace $\bm{W}_{i+1}$ for FA)
@@
  $\bm{C}=\hat{\bm{x}}_{i}\hat{\bm{x}}_{i}^{\top}$ $\triangleright$ Compute remaining correlation
  $\bm{g}=\sqrt{\bm{x}_{i}^{2}/\hat{\bm{x}}_{i}^{2}}$ $\triangleright$ Compute rescaling
  $\bm{M}_{i}\leftarrow\bm{g}\circ(\bm{M}_{i}-\eta_{M}\bm{C}\bm{M}_{i})$ $\triangleright$ Update decorrelation matrix
  $\bm{\mu}_{i}\leftarrow\bm{\mu}_{i}+0.1(\bm{x}_{i}-\bm{\mu}_{i})$ $\triangleright$ Update mean estimate (fixed learning rate)
@@
  $\bm{W}_{i+1}\leftarrow\bm{W}_{i+1}-\eta_{W}\bm{\delta}_{i+1}\hat{\bm{x}}_{i}^{\top}$ $\triangleright$ Update by SGD (alternatively by other optimizer)
@@
Appendix F Architectures and Hyperparameters